OpenStudy (anonymous):

List the first four terms of a geometric sequence with t1 = 4 and tn = -3tn-1

OpenStudy (aum):

Is this the correct problem? \(\large t_1 = 4\) \(\large t_n = -3t_{n-1}\)

OpenStudy (anonymous):

yep

OpenStudy (aum):

"n" stands for the number of the term: 1, 2, 3, etc. \(t_n\) stands for the \(n^{th}\) term. \(t_1\) is the first term, \(t_2\) is the second term, etc. So if you want the second term, put n = 2 in \(\large t_n = -3t_{n-1}\) What do you get?

OpenStudy (anonymous):

-3t2-1?

OpenStudy (aum):

\(\large t_n = -3t_{n-1}\) put n = 2 \(\large t_2 = -3t_{2-1} = -3 * t_1\) Since \(\large t_1 = 4\), \(\large t_2 = -3 * 4 = -12\) put n = 3 in \(\large t_n = -3t_{n-1}\) to find \(\large t_3\). What is the number you get for \(\large t_3\) ?

OpenStudy (aum):

A quick way to solve this problem is to interpret what \(\large t_n = -3 * t_{n-1}\) means. It says, any term is -3 times the previous term. So if first term is 4, the second term is -3 times the first term or -3 * 4 = -12. The third term is -3 times the second term = -3 * -12 = 36 The fourth term is -3 times the third term = -3 * 36 = -108 So the first four terms of the sequence are: 4, -12, 36, -108

OpenStudy (anonymous):

oh okay!! thank you so much!

OpenStudy (aum):

You are welcome.